# A Non-Expert’s Introduction to Data Ethics for Mathematicians

### v2

@@ 6.3. AI Ethics Guidelines, Tradeoffs in Different Ethical Values, and Theoretical Barriers
It is mathematically impossible for an AI system to simultaneously include many parameters, be robust to poisoning (e.g., with fake data) by an adversarial actor, and preserve privacy [${}^{+}$23], so the use of AI will always include tradeoffs between different ethical values even when everything works perfectly. Of course, AI systems don’t work perfectly. For example, it was proven recently that one can plant undetectable backdoors into machine-learning models and thereby confuse “adversarially robust” classifiers, demonstrating a major theoretical barrier to the certification of adversarial robustness [GKVZ22].

### v3

@@ 6.3. AI Ethics Guidelines, Tradeoffs in Different Ethical Values, and Theoretical Barriers
It is mathematically impossible for an AI system to simultaneously include many parameters, be robust to poisoning (e.g., with fake data) by an adversarial actor, and preserve privacy [EMFG+23], so the use of AI will always include tradeoffs between different ethical values even when everything works perfectly. Of course, AI systems don’t work perfectly. For example, it was proven recently that one can plant undetectable backdoors into machine-learning models and thereby confuse “adversarially robust” classifiers, demonstrating a major theoretical barrier to the certification of adversarial robustness [GKVZ22].